OpenStudy (anonymous):

The circumference of a circle and the perimeter of a rectangle are equal. The radius of the circle is equal to the width of the rectangle. If the radius of the circle is 6.28 ft, what is the length of the rectangle? Use 3.14 for pi

OpenStudy (lalaly):

\[2\pi r=2(l+w)\]we know that \[r=w\]so\[2\pi w=2(l+w)\]\[\pi w=l+w\]\[\pi w-w=l\]\[l=3.14 \times (6.28)-6,28\]